OpenStudy (anonymous):

f (x) = 4x g(x) = x + 5 f(x) x g(x)= 4(X) x (X) + 5 4 x 4 x 5 x 4 am i supposed to multiply or add this 16 x 20 im so stuck x = 320 but that makes no sense Help please

OpenStudy (ajprincess):

\(f(x)*g(x)=4x(x+5)=4x^2+20x\) I suppose u need to find the value of f(x)*g(x) when x=320. Am I right?

OpenStudy (anonymous):

are you asking me if thats right?

OpenStudy (anonymous):

i got 320, ut i didnt do it like you

OpenStudy (ajprincess):

what I am asking is what do u need to find?

OpenStudy (anonymous):

oh sorry i need to find f(x) x g(x)

OpenStudy (ajprincess):

it's k:) what is 4 then?

OpenStudy (anonymous):

f(x)

OpenStudy (anonymous):

right?

OpenStudy (ajprincess):

u need to find only f(x)*g(x) right? is the value of x given to u?

OpenStudy (anonymous):

right, and no

OpenStudy (ajprincess):

ok fine.:) f(x) and g(x) are given to you as 4x and x+5 respectively. To find f(x)*g(x) u jst have to multiply them. i.e f(x)*g(x)=4x(x+5)=(4x*x)+(4x*5)=4x^2+20x. getting this?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so it would be 36

OpenStudy (ajprincess):

hw did u get 36?

OpenStudy (anonymous):

4 x ^ 2 is 16x + 20 = 36 right?

OpenStudy (ajprincess):

4x^2 is not equal to 16x. 4x^2=4*x*x 16x=4*4*x getting the difference?

OpenStudy (anonymous):

oh i see so it would be asically 4 + 20

OpenStudy (anonymous):

to the power of 2

OpenStudy (ajprincess):

nope. r u familiar with like terms?

OpenStudy (anonymous):

sorry i dont get math

OpenStudy (anonymous):

yes i am

OpenStudy (ajprincess):

gud.:) Are x and 2x like terms?

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

gud:) hw abt x and 2? R they like terms?

OpenStudy (anonymous):

no

OpenStudy (ajprincess):

yup u r right:) remember u can do addition or subtraction only to like terms. If so can u add 2x and 3 and get 5?

OpenStudy (anonymous):

no you can't

OpenStudy (ajprincess):

gud:) one more question are x^2 and x like terms?

OpenStudy (anonymous):

hmmmm... this one is tough i wouldnt think so because their both x, but one is squared and one is nto, ut im still going to say yes they are like terms

OpenStudy (anonymous):

but thats wrong isnt it

OpenStudy (ajprincess):

they are nt like terms. if both the terms are in x then u need to consider the powers. the power of the first one is 2 and the second one is 1 as u said, since the powers are different they are not like terms. for like terms the powers are same.

OpenStudy (anonymous):

OH yeah that makes sense

OpenStudy (ajprincess):

now that u knw they are not like terms is it possible to add x^2 and 2x and get 3x?

OpenStudy (anonymous):

no

OpenStudy (ajprincess):

yup u r right:) is it possible to add 4x^2 and 20x and get 24?

OpenStudy (anonymous):

nope i get it now

OpenStudy (anonymous):

so i multiply?

OpenStudy (ajprincess):

no u jst leave it as it is. u need to simplify it only if it is possible.

OpenStudy (anonymous):

so now the question is it possible to be simplified?

OpenStudy (ajprincess):

yup is it possible?

OpenStudy (anonymous):

exactly, so i simplify it by

OpenStudy (ajprincess):

u mean it is possible.?

OpenStudy (anonymous):

no it no but in equations that are how do i simplify it

OpenStudy (ajprincess):

it is not possible to simplify. so u need to leave 4x^2+20x as it is. this because 4x^2 and 20x are not like terms. so u can't add them.

OpenStudy (anonymous):

i know but im saying

OpenStudy (anonymous):

in another equation how do i simplify

OpenStudy (anonymous):

if it can be simplified

OpenStudy (anonymous):

sorry im causing u much trouble but thnk you so much

OpenStudy (ajprincess):

ok say u hav to add 4x and 7x. as u can see the powers of the x in 4x and 7x are 1. so they are like terms. so u can add them as u would add numbers. 4x+7x=11x

OpenStudy (anonymous):

oh i see thnk you so much i get it now

OpenStudy (anonymous):

thnk u so much

OpenStudy (ajprincess):

u r welcome:) nw can u tell what is f(x)*g(x)=?

OpenStudy (anonymous):

yup , its 4x(x+5) = (4 x X x ) + ( 4x X 5)= 4 x^2 + 20x

OpenStudy (anonymous):

right?

OpenStudy (ajprincess):

yup right:) gud:)

OpenStudy (anonymous):

phew